OpenStudy (anonymous):

Determine whether (-3, -5) is a solution of 6x + 7y = -4

OpenStudy (happykiddo):

are there options

OpenStudy (anonymous):

no options, just either, yes or no

OpenStudy (anonymous):

Well, take -3 and plug it into x. Then take -5 and plug it into y, and if those values together equals -4, then yes it is or no it isn't.

OpenStudy (anonymous):

So it is?

OpenStudy (p0sitr0n):

6 *-3 +7*-5 = -4 -18-35=-4 obviously not